OpenStudy (anonymous):

x^2-3x+2 find all zeros of f. {-8,1,2} {-8,-1,-2} {8,-1,-2} {8,1,2}

OpenStudy (anonymous):

I don't understand, b/c 8 is not a zero, and that is in every possible answer

OpenStudy (anonymous):

this was the original question Use synthetic division to divide x3 - 11x2 + 26x - 16 by (x - 8). Use the result to find all zeros of f.

OpenStudy (anonymous):

\[0=x^2-3x+2\]Solve this by factoring or using the quadratic. Since it is a second degree poly there will be 2 answers

OpenStudy (whpalmer4):

Okay, (x-8) term contributes a zero at x=8. As you've got your possible answers already, it's easy to just try them to see which work, or use the quadratic or factor as ChmE says

OpenStudy (anonymous):

Ok, so they are giving you (x-8) as a factor already. So you know right away that 8 is a zero. So then you do the division to find the poly that contains the other 2 zeros

OpenStudy (anonymous):

i got the 1,2 im just not sure which 8 it is -_-

OpenStudy (anonymous):

1,2 is correct.

OpenStudy (anonymous):

(x-8) what is the zero of this?

OpenStudy (anonymous):

8...

OpenStudy (anonymous):

correct, so D is ur answer

OpenStudy (anonymous):

Good job!!

OpenStudy (whpalmer4):

You factor your polynomial into something that looks like (x-a)(x-b)(x-c)... = 0 Now the values of x that make any of those terms = 0 are your zeros. x=8 makes (x-8) = 0, right?

OpenStudy (whpalmer4):

In control theory, you also find poles to go with those zeros. You have some function that is a fraction like \[\frac{(x+a)(x+b)(x+c)}{(x+d)(x+e)}\] A value of x that makes x+a, x+b, or x+c = 0 is a zero. A value of x that makes x+d or x+e = 0 is a pole, so called because the value of the fraction gets big really fast as you approach a pole (like a circus tent draped over the center pole).